What is the molarity of a solution made by dissolving 2.00 g of sodium carbonate in distilled water and making it up to 250 mL with distilled water in a volumetric flask?
1
Expert's answer
2020-03-10T09:09:54-0400

250 ml=0,25 L

C=n/V;

n=m/M;

n(Na2CO3)=2/106=0,02 mol

C(Na2CO3)=0,02/0,25=0,08 mol/L